# JEPA-Reasoner: Decoupling Latent Reasoning from Token Generation

### v2

@@ Abstract
While Joint-Embedding Predictive Architecture (JEPA) has emerged as a powerful architecture for learning rich latent representations, it fundamentally lacks generative abilities. Meanwhile, current latent reasoning models remain limited by the token-by-token generation paradigm, which suffers from compounding errors and heavy context dependency. To address these limitations, we proposed JEPA-Reasoner, a novel JEPA-based architecture enhanced with generative ability for latent reasoning. We augment this architecture with a separate action-talker model, Talker, to reconstruct human-readable text from latent representations produced by the JEPA-Reasoner. Our work demonstrated that decoupling latent-space reasoning from token production enables JEPA-Reasoner to produce mixed latent vectors, laying a foundation for multi-threaded reasoning and achieving superior robustness against compounding errors in autoregressive generation.
@@ 1 Introduction
The Joint-Embedding Predictive Architecture (JEPA) (i-jepa) has demonstrated strong performance in learning semantic world representations, exhibiting a superior ability to understand the world compared with traditional end-to-end generative models that operate in pixel- or token-space. By predicting abstract representations in latent space, the JEPA architecture can filter irrelevant details and preserve essential information needed for prediction (i-jepa). Such architecture has been proven to be a viable approach for representation learning and foundation model development. Various JEPA implementations, including I-JEPA (i-jepa), V-JEPA2 (v-jepa2), and M3-JEPA (m3-jepa), have shown success across various modalities and downstream tasks.
However, JEPA models are inherently non-generative (m3-jepa) because of their objective: filling missing information in the current state (i-jepa), rather than generating new content. Furthermore, the predictor often requires a detailed, predetermined target state or task instruction (e.g., V-JEPA 2-AC (v-jepa2)), which is typically unavailable in tasks that demand long-term planning and step-by-step reasoning. This feature limits the application of the broad knowledge in JEPA models to generative tasks.
Furthermore, while traditional token-level autoregressive models have sequential reasoning capabilities, their token-by-token generation process is prone to compounding errors. Even if the predicted probabilities of tokens are weighted and combined, LLMs still cannot go beyond single-threaded reasoners (llms-reason-singel-thread). Although several previous studies have explored latent space reasoning for Transformer models, the end-to-end generation goal of a single coupled model has limited the full potential of latent space reasoning and made training complex and inefficient (coconut).
To address these limitations, we propose JEPA-Reasoner, a novel decoupled architecture that utilizes separate models for reasoning and token generation. The reasoning model, JEPA-Reasoner, transforms the JEPA framework from a target-conditioned system into an autoregressive generative model. Operating entirely within the continuous, normalized latent space, JEPA-Reasoner focuses solely on latent-space reasoning, offloading the token-generation task to its action-taker module, Talker. In contrast to existing latent space reasoning solutions like COCONUT, the division of objective frees the reasoner from expression burden and enables continuous latent guidance (compared with non-continuous latent guidance that relies on context information to retrieve latent reasoning results (coconut)) during the token generation process.
Our key insight is that performing pure reasoning in an abstract, continuous, and normalized latent space without the burden of token generation enables the construction of high-level reasoning chains that carry rich semantic information throughout the autoregressive process while correctly ignoring irrelevant details or distracting information. Our empirical experiments demonstrate further improvements beyond that: by operating in the latent space, JEPA-Reasoner can maintain multiple hypotheses simultaneously during the reasoning process and mitigate catastrophic error propagation associated with discrete token sampling.
@@ 2 Related Work
JEPA (i-jepa) introduced a framework that makes predictions in representation space. It utilizes self-supervised training to learn latent states that are not directly human-readable. A predictor module was trained to predict the target state based on encoded inputs. Multiple variants of this architecture have extended the JEPA family to various modalities and downstream tasks (i-jepa). However, these models are non-generative. Attempts to make JEPA generative, such as D-JEPA (d-jepa), use learned representations to condition diffusion models for data generation (e.g., text-to-image, text-to-audio), but they still fail to enable sequential reasoning or planning within the JEPA framework itself. In contrast, our key innovation is to adapt the core JEPA objective for autoregressive latent-space generation.
@@
Previous work on latent space reasoning mainly focuses on looping hidden states, either through horizontal autoregression like COCONUT (coconut), or vertical recurrent depth scaling (recurrent-latent-reasoning). However, these paradigms utilize a single coupled model for both latent reasoning and token generation, ignoring the mismatch between the two tasks: one requires high-level global planning, decision-making, and choice tracing, while the other requires local correctness in grammar and fluency. Besides, whether iterating over tokens or deepening the computation per token, both paradigms remain bound by causal masking, meaning early token generation cannot be guided by future reasoning states. Given these limitations, our key innovation is to decouple latent-space reasoning and token generation. By generating a complete latent chain first, we enable consistent latent guidance, with all tokens generated from the whole reasoning trajectory. This produces a higher-quality answer that is less prone to error propagation. Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models like COCONUT or Recurrent Depth transformers, which require multiple synchronized passes or complex recurrent unrolling.
@@
Modern Transformer models conduct token-level prediction in an autoregressive manner. While proven powerful on various tasks, this approach is known to suffer from compounding errors in long-horizon tasks. Techniques like Chain-of-Thought (cot) improve reasoning by generating intermediate steps, but still operate at the token level. JEPA-Reasoner aims to enhance robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
@@ 3 Model Architecture
  • Talker: Translates the latent states into tokens. Note that the Talker is not able to make predictions. Its task is to reconstruct tokens based entirely on the latent output from JEPA-Reasoner.
@@ 3.1 JEPA-Reasoner
Figure 1: Architecture of JEPA-Reasoner and its action taker (Talker). The Reasoner consists of an embedding layer as the token encoder and Transformer blocks as the predictor. The embedding layer for input tokens always uses the latest weights, while the weight of the embedding layer for target tokens is the exponential moving average of the historical weights of the input embedding layer.
@@
JEPA-Reasoner follows the JEPA philosophy, containing an embedding layer as a textual token encoder and modified Transformer blocks for the predictor, since the Transformer has proven its strong ability in sequence modeling. After the modified Transformer blocks, we applied a hybrid normalization layer (RMS and L2 normalization). We utilize L2 normalization to prevent exploding magnitude caused by residual connections. In the modified Transformer block, we apply a non-learnable QK-Norm (qk-norm) to ensure numerical stability.
@@
Unlike traditional JEPA models, in which the predictor is aimed at filling missing information in the current state (i-jepa), the predictor of JEPA-Reasoner generates the next latent matrix representing the subsequent reasoning steps. Crucially, this generated latent matrix is not projected into vocabulary probabilities via an LM head. Instead, it is normalized by the hybrid normalization layer to the unit hypersphere and looped back as the input of the first Transformer block for the next round of autoregressive generation in the latent space.
@@
The model is trained to predict latent representations provided by a target encoder. Following standard JEPA methodology, the target encoder weights are an exponential moving average (EMA) of the data encoder weights, providing stable and rich training targets. Given the normalized nature of our latent space, we use scaled cosine similarity loss computed entirely in latent space, ensuring the predictor learns consistent feature representations and dynamics (refer to Section 4.2 for more detail).
3.2 Action-Taker Model
In this experiment, there is only one action-taker model: Talker. The Talker model is a standard Transformer-based model trained independently. We designed two variants of Talker: Mono-Talker and Dual-Talker. Detailed information about the components of the two Talkers is shown in Table 1. Mono-Talker consists of decoders only, without an embedding layer or encoders. Mono-Talker is designed for reconstruction tasks that do not require context information, receiving latent vectors from JEPA-Reasoner and constructs the complete token sequence in one forward pass. Dual-Talker is designed for context-aware reconstruction, usually necessary in natural language tasks. It has an embedding layer, encoders, and decoders. The embedding layer is used for encoding previously determined outputs of JEPA-Reasoner that contain contextual information, while the encoder blocks receive latent vectors from JEPA-Reasoner as input. The decoders generate tokens autoregressively conditioned on previous tokens, with continuous latent guidance from the output of encoders. However, Dual-Talker was trained for reconstruction rather than generation, as our ablation study (Appendix C) showed that Talker is critically dependent on the Reasoner’s output. During training, the JEPA-Reasoner is frozen. Talker receives the sequence of latent vectors and is trained to reconstruct the corresponding token sequence using standard cross-entropy loss.
@@ 4.2 Self-Supervised Training
The SST phase transforms the model from a token generator into a pure latent-space reasoner. We adopt a self-supervised approach inspired by the JEPA series (i-jepa). Unlike methods such as COCONUT (coconut), our approach avoids the computational overhead of autoregressively generating final tokens or looping latent vectors multiple times for loss computation, enabling more efficient parallel training.
@@
The model is now optimized to predict the latent representation of the next sequence segment. We discard the temporary LM head and restore the L2 normalization layer. We apply a similar self-supervised training strategy proposed in Meta’s JEPA series (i-jepa) and switch to scaled cosine distance loss, aligning with the L2 normalization strategy to ensure stability during autoregressive looping and encouraging the model to learn angular differences:
@@
where $k$ is a scalar, $h_{\text{pred}}$ is the latent vector predicted by the Reasoner (parameters $\theta$) and $h_{\text{target}}$ is the target vector from the EMA encoder (parameters $\theta^{\prime}$). We adopt a scaling factor of $k=4$ as we empirically observed that standard cosine distance yields insufficient gradients when the loss is small. (See Appendix D for more discussions about scaled cosine distance loss).
@@ 5.1.4 Results and Conclusions
We calculated the distance from the predicted latent vector to the plane spanned by any two vocabulary vectors and sorted them from closest to farthest. In the sorted list, the plane spanned by latent vectors of sibling nodes frequently exhibits lower distances to the predicted latent vector, with an average ranking of top 1.72% in the ordered list. Also, we figured out all coefficient sets, $\alpha$ and $\beta$, that satisfy $\alpha\cdot\mathbf{l}_{0}+\beta\cdot\mathbf{l}_{1}=\mathbf{l}_{proj}$, where $\mathbf{l}_{0}$ and $\mathbf{l}_{0}$ are latent vectors of sibling nodes and $\mathbf{l}_{proj}$ is the projection of the predicted latent vector on the spanned plane. After comparing $\alpha$ and $\beta$, we find that for 99.89% of the time, the latent vector of the node on the correct route contributes more than the other sibling node. This discovery demonstrated that JEPA-Reasoner could make correct choices without completely discarding the other information that contains potentially correct choices. According to the previous COCONUT study (coconut), this behavior might lay the foundation for breadth-first multi-threaded reasoning.
@@ 5.2 Robustness to Error Perturbation
We demonstrate that decoupling reasoning generation from token production yields superior robustness. While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on logical consistency in the latent space, demonstrating superior generation quality and robustness to noisy input. We validate this advantage through two experiments targeting token-level errors and latent space noise, respectively.
@@ 5.2.1 Experiment Methods
To evaluate the robustness of the decoupled model on token-level errors in the input sequence, we randomly replace 0% to 30% ground truth tokens in the input sequence with incorrect tokens. We compare the performance of JEPA-Reasoner and traditional Transformer models on multi-step completion tasks [^1] All scores obtained in these two tests are by testing the model across 5248 samples randomly chosen from the test dataset containing 100000 samples to minimize the bias introduced by randomness. using the exact match metric.
@@
To evaluate the robustness of the JEPA-Reasoner model on perturbations in latent space, we compare the performance of JEPA-Reasoner and the coupled continuous reasoning model COCONUT on multi-step completion tasks. In this experiment, COCONUT generates four latent vectors followed by four tokens, whereas JEPA-Reasoner generates eight latent vectors and reconstructs eight tokens via the Talker module. We inject Gaussian noise at each step ($\mu=0$; $\sigma\in[0\%,15\%]$ relative to the maximum output value) and measure exact match accuracy on the final four tokens.
@@ 5.2.2 Data Preparation
Considering that the compounding error caused by different faulty tokens differs significantly, it is difficult to quantitatively analyze the model’s behavior under token-level errors (e.g., replacing keywords in the sentence will decrease the quality more considerably than replacing a word that functions as a connector). We followed previous work (allen-physics) and utilized Context-Free Grammar (CFG) for both experiments to create a controllable experiment setting.
Our custom CFG production rule features three terminal symbols with rule lengths of 3 or 4. With this rule, we generated long (approximately 600 to 700 symbols) and complex sequences that require non-trivial work to solve. The complexity of the grammar ensures that high accuracy relies on learning the underlying structure of the CFG sequence rather than memorizing specific sequences (refer to Appendix B.1 for full CFG specifications and production methods).

### v3

@@ Abstract
Current autoregressive language models couple high-level reasoning and low-level token generation into a single sequential process, making the reasoning trajectory vulnerable to compounding expression errors. We propose JEPA-Reasoner, a novel architectural paradigm that decouples these tasks using a Joint-Embedding Predictive Architecture (JEPA) for pure latent-space reasoning and a separate Talker module for linguistic reconstruction. By isolating the reasoning engine from the discrete token-sampling process, our architecture enables: (1) Error Containment, where token-level failures cannot propagate into the latent reasoning chain; (2) Continuous Guidance, providing the generator with access to the entire lossless reasoning trajectory; and (3) Representation of Uncertainty, allowing the model to maintain multiple hypotheses via mixed latent vectors. Controlled experiments on synthetic and natural language tasks demonstrate that this decoupling enables a 0.9B model to achieve a 149.5% improvement in 8-shot GSM8K accuracy over a coupled Transformer baseline trained on identical data. This work shifts the focus from scaling coupled models to investigating decoupled architectures as a more robust foundation for complex reasoning.
@@ 1 Introduction
The dominant paradigm for Large Language Models (LLMs) is a unified autoregressive process where internal reasoning and external expression are inextricably linked. In this coupled framework, the model must predict the next token based on a history that includes its own previously sampled discrete tokens. This creates a fundamental bottleneck: any localized error in token selection—whether a grammatical slip or a sub-optimal word choice—pollutes the context window, directly corrupting the trajectory of all subsequent reasoning.
While recent efforts such as COCONUT (Hao et al., 2024) and Recurrent Depth models (Geiping et al., 2025) have explored latent-space reasoning, they remain largely constrained by this coupling. They either interleave latent vectors with tokens in a single sequence or rely on causal masking that prevents early tokens from being informed by future reasoning states. Furthermore, these models struggle to represent uncertainty, as the discrete nature of token-level coupling often forces a collapse into a single reasoning path (Wu et al., 2025).
In this paper, we propose a novel decoupled reasoning paradigm: JEPA-Reasoner, an architecture that fully decouples latent-space reasoning from token generation. Our approach utilizes the Joint-Embedding Predictive Architecture (JEPA) (Assran et al., 2023) not merely for representation learning, but as an independent, autoregressive reasoning engine. The Reasoner operates entirely within a continuous, normalized latent space, generating a complete reasoning chain before any tokens are produced. A separate, lightweight Talker module then reconstructs human-readable text under the guidance of this completed latent trajectory.
By isolating these functions, we demonstrate three emergent properties of the decoupled architecture:
  1. 1.
    Mathematical Error Containment: We provide a probabilistic factorization showing that because the reasoning chain is generated independently, token-sampling errors have no mathematical pathway to influence the logic of the Reasoner.
  2. 2.
    Architectural Efficiency: Our experiments show that JEPA-Reasoner utilizes internally learned knowledge more effectively than coupled baselines. When trained on identical datasets, the decoupled architecture overcomes the ”reasoning plateau” observed in small-scale models, showing superior in-context scaling from 5-shot to 8-shot tasks.
  3. 3.
    Mixed Latent States: We empirically validate that the Reasoner can produce ”mixed latent vectors” that encode multiple possible reasoning paths simultaneously, laying the foundation for multi-threaded reasoning without the overhead of beam search or multiple decodes.
Rather than focusing on competitive SOTA benchmarks through massive scaling, this work serves as an architectural investigation into the benefits of decoupling. We show that even without sophisticated reinforcement learning or industrial training recipes, transitioning from a coupled to a decoupled paradigm yields significant gains in robustness and reasoning depth, suggesting a new path forward for foundation model design.
@@ 2 Related Work
JEPA (Assran et al., 2023) introduced a framework that makes predictions in representation space. It utilizes self-supervised training to learn latent states that are not directly human-readable. A predictor module was trained to predict the target state based on encoded inputs. Multiple variants of this architecture have extended the JEPA family to various modalities and downstream tasks (Assran et al., 2023). However, these models are non-generative. Attempts to make JEPA generative, such as D-JEPA (Chen et al., 2025), use learned representations to condition diffusion models for data generation (e.g., text-to-image, text-to-audio), still fail to enable sequential reasoning or planning within the JEPA framework itself. In contrast, our key innovation is to adapt the core JEPA objective for autoregressive latent-space generation.
@@
Previous work on latent space reasoning mainly focuses on looping hidden states, either through horizontal autoregression like COCONUT (Hao et al., 2024), or vertical recurrent depth scaling (Geiping et al., 2025). However, these paradigms utilize a single coupled model for both latent reasoning and token generation, ignoring the mismatch between the two tasks: one requires high-level global planning, decision-making, and choice tracing, while the other requires local correctness in grammar and fluency. Besides, whether iterating over tokens or deepening the computation per token, both paradigms remain bound by causal masking, meaning early token generation cannot be guided by future reasoning states. Given these limitations, our key innovation is to decouple latent-space reasoning and token generation. By generating a complete latent chain first, we enable consistent latent guidance, with all tokens generated from the whole reasoning trajectory. This produces a higher-quality answer that is less prone to error propagation. Additionally, the decoupled design enables efficient optimization with a single forward pass in latent space, unlike coupled models such as COCONUT or Recurrent Depth Transformers, which require multiple synchronized passes or complex recurrent unrolling.
@@
Modern Transformer models conduct token-level prediction in an autoregressive manner. While proven powerful on various tasks, this approach is known to suffer from compounding errors in long-horizon tasks. Techniques like Chain-of-Thought (Wei et al., 2022) improve reasoning by generating intermediate steps, but still operate at the token level. JEPA-Reasoner aims to enhance robustness of autoregressive generation by moving the reasoning process into a continuous, abstract latent space, reducing the impact of localized errors.
@@ 3 Model Architecture
  • Talker: Translates the latent states into tokens. Note that the Talker can not make predictions. (Refer to ablation study in Appendix C) Its task is to reconstruct tokens solely based on the latent reasoning chain produced by JEPA-Reasoner.
@@ 3.1 JEPA-Reasoner
Figure 1: Architecture of JEPA-Reasoner and Talker. The Reasoner consists of an embedding layer as the token encoder and Transformer blocks as the predictor. The embedding layer for input tokens always uses the latest weights, while the weight of the embedding layer for target tokens is the exponential moving average of the historical weights of the input embedding layer.
@@
JEPA-Reasoner follows the JEPA philosophy, containing an embedding layer as a textual token encoder and modified Transformer blocks for the predictor, since the Transformer has proven its strong ability in sequence modeling. After the modified Transformer blocks, we applied a hybrid normalization layer (RMS and L2 normalization). We utilize L2 normalization to prevent exploding magnitude caused by residual connections. In the modified Transformer block, we apply a non-learnable QK-Norm (Dehghani et al., 2023) to ensure numerical stability.
@@
Unlike traditional JEPA models, in which the predictor is aimed at filling missing information in the current state (Assran et al., 2023), the predictor of JEPA-Reasoner generates the next latent matrix representing the subsequent reasoning steps. Crucially, this generated latent matrix is not projected into vocabulary probabilities via an LM head. Instead, it is normalized by the hybrid normalization layer to the unit hypersphere and looped back as the input of the first Transformer block for the next round of autoregressive generation in the latent space.
@@
The model is trained to predict latent representations provided by a target encoder. Following standard JEPA methodology, the target encoder weights are an exponential moving average (EMA) of the data encoder weights, providing stable and rich training targets. Given the normalized nature of our latent space, we use scaled cosine similarity loss computed entirely in the latent space, ensuring the predictor learns consistent feature representations and dynamics (refer to Section 4.2 for more details).
3.2 Talker Model
The Talker model is a standard Transformer-based model trained independently. We designed two variants of Talker: Mono-Talker and Dual-Talker. Detailed information about the components of the two Talkers is shown in Table 1. Mono-Talker consists of decoders only, without an embedding layer or encoders. Mono-Talker is designed for reconstruction tasks that do not require context information, receiving latent vectors from JEPA-Reasoner and constructs the complete token sequence in one forward pass. Dual-Talker is designed for context-aware reconstruction, usually necessary in natural language tasks. It has an embedding layer, encoders, and decoders. The embedding layer is used for encoding previously determined outputs of JEPA-Reasoner that contain contextual information, while the encoder blocks receive latent vectors from JEPA-Reasoner as input. The decoders generate tokens autoregressively conditioned on previous tokens, with continuous latent guidance from the output of encoders. However, Dual-Talker was trained for reconstruction rather than generation, as our ablation study (Appendix C) showed that Talker is critically dependent on the Reasoner’s output. During training, the JEPA-Reasoner is frozen. Talker receives the sequence of latent vectors and is trained to reconstruct the corresponding token sequence using standard cross-entropy loss.
@@ 4.2 Self-Supervised Training
The SST phase transforms the model from a token generator into a pure latent-space reasoner. We adopt a self-supervised approach inspired by the JEPA series (Assran et al., 2023). Unlike methods such as COCONUT (Hao et al., 2024), our approach avoids the computational overhead of autoregressively generating final tokens or looping latent vectors multiple times for loss computation, enabling more efficient parallel training.
@@
The model is now optimized to predict the latent representation of the next sequence segment. We discard the temporary LM head and restore the L2 normalization layer. We apply a similar self-supervised training strategy proposed in Meta’s JEPA series (Assran et al., 2023) and switch to scaled cosine distance loss, aligning with the L2 normalization strategy to ensure stability during autoregressive looping and encouraging the model to learn angular differences:
@@
where $k$ is a scalar, $h_{\text{pred}}$ is the latent vector predicted by the Reasoner (parameters $\theta$) and $h_{\text{target}}$ is the target vector from the EMA encoder (parameters $\theta^{\prime}$). We adopt a scaling factor of $k=4$ as we empirically observed that standard cosine distance yields insufficient gradients when the loss is small. (See Appendix D for detailed discussions about scaled cosine distance loss).
@@ 5.1.4 Results and Conclusions
Visualization of Mixed Latent Vectors
To visualize the mixed latent vectors, we extracted embedding vectors from distinct tree leaves alongside the model’s output latent representations after one forward pass. Principal Component Analysis (PCA) was applied to the collected embeddings and model predictions, and Figure 2 demonstrates the visualization of first two principal components (PC1 and PC2). Noticing that predicted latent vectors (blue points in the figure) form a continuous cloud within the space spanned by discrete vocabulary embeddings (red diamond shapes). This distribution supports the experiment results that they are the linear combinations of vocabulary embeddings. Also, the predicted vectors do not converge to singular vocabulary points, providing empirical evidence for the hypothesis that JEPA-Reasoner is capable of maintaining information from multiple possible choices rather than committing to a single answer.
Quantitative Analysis
We calculated the distance from the predicted latent vector to the plane spanned by any two vocabulary vectors and sorted them from closest to farthest. In the sorted list, the plane spanned by latent vectors of sibling nodes frequently exhibits lower distances to the predicted latent vector, with an average ranking of top 1.72% in the ordered list. Also, we figured out all coefficient sets, $\alpha$ and $\beta$, that satisfy $\alpha\cdot\mathbf{l}_{0}+\beta\cdot\mathbf{l}_{1}=\mathbf{l}_{proj}$, where $\mathbf{l}_{0}$ and $\mathbf{l}_{0}$ are latent vectors of sibling nodes and $\mathbf{l}_{proj}$ is the projection of the predicted latent vector on the spanned plane. After comparing $\alpha$ and $\beta$, we find that for 99.89% of the time, the latent vector of the node on the correct route contributes more than the other sibling node. This discovery demonstrated that JEPA-Reasoner could make correct choices without completely discarding the other information that contains potentially correct choices. According to the previous COCONUT study (Hao et al., 2024), this behavior might lay the foundation for breadth-first multi-threaded reasoning.
@@ 5.2 Robustness to Error Perturbation
We demonstrate that decoupling reasoning generation from token production yields superior robustness. While coupled models must simultaneously maintain reasoning coherence and produce correct tokens, our decoupled approach allows the reasoning model to focus solely on consistency in the latent space, demonstrating superior generation quality and robustness to noisy input. We validate this advantage through two experiments targeting token-level errors and latent space noise, respectively.
@@ 5.2.1 Experiment Methods
To evaluate the robustness of the decoupled model on token-level errors in the input sequence, we randomly replace 0% to 30% ground truth tokens in the input sequence with incorrect tokens. We compare the JEPA-Reasoner against traditional Transformer models on multi-step completion tasks using the exact match metric. To minimize bias, we test on 5,248 samples randomly selected from a test set of 100,000.
@@
To evaluate JEPA-Reasoner’s robustness against perturbations in latent space, we compare its performance with the coupled continuous reasoning model COCONUT on multi-step completion tasks. In this setup, COCONUT generates four latent vectors followed by four tokens, whereas JEPA-Reasoner generates eight latent vectors and reconstructs eight tokens via the Talker module. We inject Gaussian noise at each step ($\mu=0$; $\sigma\in[0\%,15\%]$ relative to the maximum value in each model’s output latent vectors) and measure exact match accuracy on the final four tokens.
@@ 5.2.2 Data Preparation
Quantifying and testing token-level robustness in natural language is challenging because the impact of error tokens varies significantly by token type (e.g., replacing a keyword is more damaging than replacing a connector). To create a controlled experiment setup, we followed previous work (Allen-Zhu and Li, 2023) and utilized Context-Free Grammar (CFG) for both experiments.
Our custom CFG features three terminal symbols with rule lengths of 3 or 4, producing complex and long sequences of 600–700 symbols. The diversity of sequences generated ensures that high accuracy requires learning the underlying structure rather than simple memorization (see Appendix B for detailed discussion).